OpenStudy (anonymous):

I'm working with functions. If my function is g(x) = (x+2)^2 then what is g(i-3)?

OpenStudy (anonymous):

i^2-2i+1

OpenStudy (anonymous):

wait how did you get that?

OpenStudy (anonymous):

>.> bump

OpenStudy (saifoo.khan):

im here!

OpenStudy (anonymous):

hai

OpenStudy (anonymous):

so this dude just gave me the answer.. unhelpful

OpenStudy (saifoo.khan):

g(x) = (x+2)^2 insert this value: g(i-3) instead of x g(i-3) = (i-3 + 2)^2 g(i-3) = (i -1)^2 g(i-3) = (i-1)(i-1) g(i-3) = - 2 i answer^

OpenStudy (saifoo.khan):

his answer is wrong.

OpenStudy (anonymous):

OH thank you :P Geeze I'm blanking. This website is so cool :)

OpenStudy (saifoo.khan):

thnks! ^.^

OpenStudy (anonymous):

I have another question. How is it - 2 i? wouldn't you get like i^2 -2i +1?

OpenStudy (saifoo.khan):

coz i^2 is -1 not i^2

OpenStudy (anonymous):

Geeze I knew that and I couldn't figure it out haha.